OpenStudy (anonymous):

Solve by elimination. 5x + y = 5 9x - 4y = -20

OpenStudy (disco619):

Ok, i don't really know If Im doing this the right way but here I go: 5x + y = 5 9x - 4y = -20 5x + y = 5 => x4 9x - 4y = -20 20x + 4y = 20 9x - 4y = -20 29x = 0 20x/29 = 0/29 x = 0 LOL I think this might be Wrong. @mohammad.arqum You tell...

OpenStudy (anonymous):

ok...its oke

OpenStudy (disco619):

I guess it's right then! :D

OpenStudy (anonymous):

hmmm well the problem is multiple choice. The choices are: A. (15/29, -220/29) B. (-15/29, 220/29) C. (0,-5) D. (0,5)

OpenStudy (anonymous):

ok when x=0 then put the value of x in the first equation... 5x+y=5 5(0) + y = 5 0 + y = 5 y=5 therefore the solution is (0,5) option D

OpenStudy (disco619):

The answer is "D"

OpenStudy (anonymous):

Thank you both!!!!! (:

OpenStudy (disco619):

You're Welcome! Anytime... :)

OpenStudy (only4u):

yes (0,5) is the answer